OpenStudy (anonymous):

CAN SOMEONE HELP ME Which of the following choices is the simple formula for the nth term of the following arithmetic sequence? 5, 1, -3, -7, -11, ... A. -4n + 1 B. 4n + 1 C. -4n + 9 D. 5n + 4 E. 4n + 9

OpenStudy (anonymous):

The common difference between terms is the factor of n. What is the common difference in your question?

OpenStudy (anonymous):

"multiplier of n", Sorry

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Not quite. The common difference is the difference between any term and the term immediately before it. In other words, what do you have to add to 5 in order to get 1?

OpenStudy (anonymous):

4-5=1

OpenStudy (anonymous):

o

OpenStudy (anonymous):

If we use d to represent common difference, then\[d=n _{k}-n _{k-1}\]In the given question, d=1-5. What do you get?

OpenStudy (anonymous):

1-5=4

OpenStudy (anonymous):

so 4 right???

OpenStudy (anonymous):

Not quite. Shouldn't it be -4?

OpenStudy (anonymous):

yea my bad lol

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

OK. So the multiplier of n has to be -4. That narrows your choices down to two.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

For the first term, n=1. So test the choices with n=1 and see which one gives you the first term, 5.

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

Right on. Well done!